Solve using the square root property. 4(x + 2)2 = 11
Answer #1

* Given equation is 4(X+21? = 11. taking square root on both sidee 4(af2)2 =UT = 2(x+2) = EVIT => af 2 fVT 2 x = £VUT 2 2 x=